# DeepAAT: Deep Automated Aerial Triangulation for Fast UAV-based Mapping

### 02 Feb 2024 v1

@@ Abstract
Automated Aerial Triangulation (AAT), aiming to restore image pose and reconstruct sparse points simultaneously, plays a pivotal role in earth observation. With its rich research heritage spanning several decades in photogrammetry, AAT has evolved into a fundamental process widely applied in large-scale Unmanned Aerial Vehicle (UAV) based mapping. Despite its advancements, classic AAT methods still face challenges like low efficiency and limited robustness. This paper introduces DeepAAT, a deep learning network designed specifically for AAT of UAV imagery. DeepAAT considers both spatial and spectral characteristics of imagery, enhancing its capability to resolve erroneous matching pairs and accurately predict image poses. DeepAAT marks a significant leap in AAT’s efficiency, ensuring thorough scene coverage and precision. Its processing speed outpaces incremental AAT methods by hundreds of times and global AAT methods by tens of times while maintaining a comparable level of reconstruction accuracy. Additionally, DeepAAT’s scene clustering and merging strategy facilitate rapid localization and pose determination for large-scale UAV images, even under constrained computing resources. The experimental results demonstrate DeepAAT’s substantial improvements over conventional AAT methods, highlighting its potential in the efficiency and accuracy of UAV-based 3D reconstruction tasks. To benefit the photogrammetry society, the code of DeepAAT will be released at: https://github.com/WHU-USI3DV/DeepAAT.
Index Terms:
Automated Aerial Triangulation (AAT), Unmanned Aerial Vehicle (UAV), Structure from Motion (SfM), Orientation
I Introduction
@@
Automated Aerial Triangulation (AAT) is a basic task in photogrammetry and holds substantial research significance [1]. It serves as the initial step in the 3D reconstruction pipeline of aerial images [2]. AAT’s primary role involves simultaneously recovering the camera poses and reconstructing sparse 3D points in the scene. These foundational outputs facilitate subsequent dense image matching and 3D modeling procedures [3]. The derived camera poses and scene models find diverse applications in digital mapping [4], virtual reality [5], and smart city [6]. With a research history spanning decades [7], classical AAT algorithms can be primarily categorized into two groups: incremental style and global style [8]. Furthermore, the evolution of deep learning has given rise to numerous supervised AAT algorithms [9]. The existing AAT methods are reviewed as follows.
I-A Classic Automated Aerial Triangulation
@@
The first step of the classic AAT is to perform feature extraction and matching of all input images. The following steps are different for the global style and incremental style. The global AAT can predict all camera poses and scene structure at once [10]. In AAT algorithms, Bundle Adjustment (BA) [11] is the most time-consuming part. Global AAT only requires to execute BA once, resulting in higher efficiency. However, it can be difficult to eliminate outliers, resulting in poor robustness and scene integrity. Incremental AAT was first proposed by Snavely et al. [12], whose key lies in selecting a good initial matching image pair [13]. Afterward, incremental AAT adds a new image to the system sequentially, followed by Perspective-n-Points (PnP) [14], Triangulation [15], and local BA. Incremental AAT requires multiple BA, resulting in low reconstruction efficiency in situations with a large number of images [16]. In addition, due to the accumulation of errors, the reconstructed scene is prone to drift issues.
@@
Compared to the general scenes, UAV images exhibit distinctive characteristics, including large volumes, high resolutions, and significant overlap. Within the realm of classic AAT algorithms, incremental methods have emerged as the standard approach for UAV image AAT due to their superior robustness against outliers and ability to provide comprehensive results. Addressing the challenges posed by large-scale UAV images, most SOTA AAT methods typically involve employing a divide-and-conquer strategy. This strategy begins by segmenting the UAV image into blocks based on GPS information, followed by the fusion of all modules to yield globally consistent large-scale results. Noteworthy contributions in this field include the work by Chen et al. [17], which employed the maximum spanning tree to expand images after dividing the scene map into smaller segments with a certain degree of overlap, thereby enhancing connectivity and scene map integrity. Similarly, Xu et al. [18] introduced a hierarchical approach that constructed a binary tree using images as leaf nodes, subsequently fusing spatial triads and scenes from the bottom up. This method offers advantages in terms of robustness, accuracy, and efficiency. Likewise, Bhowmick et al. [19] initially organized images into hierarchical trees using clustering methods, then addressed the AAT problem for large-scale images by reconstructing each small image set and merging them into a common reference framework. Snavely et al. [20] partitioned extensive scenes by calculating the small bone skeleton set and reconstructing the skeleton set of the image. This approach reduces the number of parameters under consideration and enhances reconstruction efficiency. To sum up, global AAT offers high efficiency but suffers from poor robustness and scene integrity. On the other hand, incremental AAT exhibits high robustness and accuracy but tends to have relatively lower time efficiency.
I-B Supervised Automated Aerial Triangulation
@@
Recognizing the limitations encountered by classical AAT algorithms, an increasing number of studies are exploring the application of deep learning methods to address these challenges. Many existing deep learning methods directly regress the depth map and pose of a monocular camera [21], which usually highly rely on prior information for prediction. In addition, due to not considering the correlation between depth and pose, the generalization ability of these networks is limited, making it difficult to obtain ideal prediction results. BA-Net [22] attempts to use feature metric BA to solve the AAT problem. It makes end-to-end training possible by designing a differentiable LM (Levenberg–Marquardt) optimization algorithm, but the LM algorithm occupies a large amount of memory and has low computational efficiency. DeepSfM [23] can simultaneously regress the pose and depth maps corresponding to the image, but it requires coarse poses and depth maps as initialization, and has high requirements for GPU, making it difficult to expand to high-resolution images and large-scale environments. DeepMLE [9] does not require initial values as input, which expresses the two-view AAT problem as maximum likelihood estimation, learning the relative pose of the two views by maximizing the likelihood function of correlation. Similarly, for the problem of binocular vision estimation, Wang et al. [24] proposed a dense optical flow estimation network for predicting between two frames, which includes a scale-invariant depth estimation module that can simultaneously calculate the relative pose of the camera based on the corresponding relationship of 2D optical flow. DRO [25] is an optimization method based on recurrent neural networks that iteratively updates camera pose and image depth to minimize feature measurement errors. Zhuang and Chandraker [26] used a self-attention graph neural network to enhance strong interactions between different corresponding relationships and potentially model complex relationships between points to drive learning. MOAC [27] introduces a grouped dual cost enhancement module, which enhances the spatial semantic information and channel relationships of costs, making the optimization more robust to noise. Moran et al. [28] proposed a new approach to solve AAT problems using deep learning. They use matched feature points as input, and after permutation equivariant networks, predict the pose of each camera and 3D points in the scene. Compared to many existing deep learning AAT methods, it can be applied to large-scale reconstruction tasks in an unsupervised manner. However, there are mainly two drawbacks to it. The first one is that it cannot eliminate incorrectly matched point pairs, which means all the input matched pairs should be correct and is usually difficult to achieve. Another one is that its prediction results are still not satisfactory because of the limited generalizability.
@@
(1) DeepAAT incorporates a spatial-spectral feature aggregation module, specifically assimilating both the spatial layout and spectral characteristics of an image set. This module boosts the network’s ability to perceive the spatial arrangement of cameras and enhances the global regression capability for poses.
(2) DeepAAT introduces an outlier rejection module rooted in global consistency, which effectively generates a reliability evaluation score for each feature correspondence. This approach facilitates the efficient and precise elimination of erroneous matching pairs, thereby ensuring accuracy and reliability throughout the entire 3D reconstruction process.
@@
The rest of this paper is structured as follows. The preliminaries for our system are provided in Section II. A brief system overview including the hardware and software structure is provided in Section III. A detailed description of DeepAAT is presented in Section IV and experiments are conducted on UAV image datasets in Section V. Conclusion and future work are drawn in Section VI.
II Preliminary
II-A Problem Definition of Automated Aerial Triangulation
The task of AAT refers to estimating the camera poses and 3D feature points corresponding to the 2D observations on the images. In classic photogrammetry, it is well studied and understood that the relative camera poses and 3D feature points can be solved only with the 2D observations [29]. The absolute camera poses and 3D feature points respective to the geodesic framework can be then obtained with Ground Control Points (GCPs) or the GPS mounted on the UAV [30].
Assume that the stationary targeting survey area is viewed by $M$ images, which are captured by the camera with known pre-calibrated intrinsic parameter $\mathbf{K}$ at different places along the UAV survey mission. The $M$ unknown camera poses are represented by a set of projection matrices $\mathcal{P}=\{\mathbf{P}_{m}|m=1,...,M\}$. Each projection matrix $\mathbf{P}_{m}$ with the size of $3\times 4$ is constructed by camera rotation $\mathbf{R}_{m}\in SO(3)$ (corresponding quaternion is $\mathbf{q}_{m}$) and position $\mathbf{t}_{m}\in\mathbb{R}^{3}$ according to $\mathbf{P}_{m}=[\mathbf{R}_{m}|\mathbf{t}_{m}]$. Given $N$ 3D feature points in the targeting survey area =$\mathcal{F}=\{\mathbf{F}_{n}|n=1,...,N\}$, each 3D feature point is written by $\mathbf{F}_{n}=[\mathbf{F}^{1}_{n},\mathbf{F}^{2}_{n},\mathbf{F}^{3}_{n},1]^{\top}$ in homogeneous coordinates. If $\mathbf{F}_{n}$ can be observed by the $m^{th}$ image, its projection on the $m$’s image is given by Eq.(1). As the depth information $\lambda_{m,n}$ is lost during the projection, $\mathbf{f}_{m,n}$ is an up-to-scale bearing vector.
@@ II-A Problem Definition of Automated Aerial Triangulation
In a typical AAT procedure, the initial step involves 2D feature detection and matching between pairwise images, which is carried out using the widely used Scale-Invariant Feature Transform (SIFT) [31] or other robust feature detectors and descriptors [32]. This step is not the main focus of this work. Subsequently, a set of 2D feature tracks denoted as $\mathcal{T}=\{\mathbf{T}_{n}|n=1,...,N\}$, is used as the input for our algorithm. It should be noted that track $\mathbf{T}_{n}$ corresponds to the 3D feature $\mathbf{F}_{n}$ and is constructed by a set of 2D observations from different images using Eq.(2):
@@
where $\mathcal{J}_{n}$ denotes the set of images that can observe the 3D feature $\mathbf{F}_{n}$. Then the tracks could be used to recover the camera poses and 3D feature points like the existing incremental [8] or global [33] AAT strategies.
II-B Projective Factorization
Despite the mainstream AAT methods listed above, projective factorization [34] is also a long-established method in AAT. We provide a brief introduction to projective factorization, as it forms the foundation for the operation of the proposed network. The complete image projections, namely the 2D feature tracks $\mathcal{T}$, can be gathered into a projective matrix $\mathbf{W}_{\mathbf{prj}}$ in Eq.(3):
$\begin{split}\mathbf{W}_{\mathbf{prj}}&\equiv\left[\begin{matrix}\lambda_{1,1}\mathbf{f}_{1,1}&\lambda_{1,2}\mathbf{f}_{1,2}&\cdots&\lambda_{1,N}\mathbf{f}_{1,N}\\ \lambda_{2,1}\mathbf{f}_{2,1}&\lambda_{2,2}\mathbf{f}_{2,2}&\cdots&\lambda_{2,N}\mathbf{f}_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{M,1}\mathbf{f}_{M,1}&\lambda_{M,2}\mathbf{f}_{M,2}&\cdots&\lambda_{M,N}\mathbf{f}_{M,N}\\ \end{matrix}\right]\\ &=\left[\begin{array}[]{c}\mathbf{KP}_{1}\\ \mathbf{KP}_{2}\\ \vdots\\ \mathbf{KP}_{M}\\ \end{array}\right]\left[\begin{array}[]{c}\mathbf{F}_{1}\\ \mathbf{F}_{2}\\ \vdots\\ \mathbf{F}_{N}\\ \end{array}\right]^{\top}.\end{split}$ (3)
If the 3D feature points in the targeting survey area $\mathcal{F}$ are observed by all the images, the camera poses and 3D feature points can be recovered using Singular Value Decomposition (SVD) of $\mathbf{W}_{\mathbf{prj}}$ [34]. For the common cases of missing observations, the SVD can be replaced with iterative methods [35, 36]. However, these methods are usually too weak for AAT in the presence of outliers and noise [37], and can not be directly applied to large-scale AAT. Nevertheless, the formulation of $\mathbf{W}_{\mathbf{prj}}$ provides an ideal way of keeping spatial correlation information for neural networks.
II-C Permutation Equivariant Layer
Let $\mathbf{W}$ be a tensor with the shape of $M\times N\times D$, whose row represents the image index, the column represents the feature track index and the third dimension represents the feature index. Taking projective matrix $\mathbf{W}_{\mathbf{prj}}$ in Eq.(3) as an example, $\mathbf{W}_{\mathbf{prj}}$ can be rearranged with the shape of $M\times N\times 2$ (the third dimension records the feature coordinates on the image plane) to serve as input for the neural network for the sake of convenience. To recover the camera poses and 3D feature points using a deep neural network, we expect a particular layer can output the same results irrespective of the order of the camera poses or the feature tracks. This reordering problem could be solved using the Permutation Equivariant Layer (PEL)[38], which was first introduced by Moran et al. [28] to handle SfM problem exploring tensor’s exchangeability.

### 07 Apr 2024 v2

@@ Abstract
Automated Aerial Triangulation (AAT), aiming to restore image poses and reconstruct sparse points simultaneously, plays a pivotal role in earth observation. AAT has evolved into a fundamental process widely applied in large-scale Unmanned Aerial Vehicle (UAV) based mapping. However classic AAT methods still face challenges like low efficiency and limited robustness. This paper introduces DeepAAT, a deep learning network designed specifically for AAT of UAV imagery. DeepAAT considers both spatial and spectral characteristics of imagery, enhancing its capability to resolve erroneous matching pairs and accurately predict image poses. DeepAAT marks a significant leap in AAT’s efficiency, ensuring thorough scene coverage and precision. Its processing speed outpaces incremental AAT methods by hundreds of times and global AAT methods by tens of times while maintaining a comparable level of reconstruction accuracy. Additionally, DeepAAT’s scene clustering and merging strategy facilitate rapid localization and pose determination for large-scale UAV images, even under constrained computing resources. The experimental results demonstrate that DeepAAT substantially improves over conventional AAT methods, highlighting its potential for increased efficiency and accuracy in UAV-based 3D reconstruction tasks. To benefit the photogrammetry society, the code of DeepAAT will be released at: https://github.com/WHU-USI3DV/DeepAAT.
keywords:
[…]
{highlights}
Incorporating a spatial-spectral feature aggregation module, boosts the network’s ability to perceive the spatial distribution of cameras and enhances the global regression capability for camera poses.
Introducing an outlier rejection module according to global consistency, which effectively generates a reliability evaluation score for each feature correspondence.
DeepAAT can efficiently process hundreds of UAV images simultaneously, marking a significant breakthrough in enhancing the applicability of deep learning-based AAT algorithms.
1 Introduction
@@
Automated Aerial Triangulation (AAT) is a basic task in photogrammetry and holds substantial research significance (Tanathong and Lee, 2014). It serves as the initial step in the 3D reconstruction pipeline of aerial images (Zhong et al., 2023). AAT’s primary role involves simultaneously recovering the camera poses and reconstructing sparse 3D points in the scene. These foundational outputs facilitate subsequent dense image matching and 3D modeling procedures Jiang et al. (2021). The derived camera poses and scene models find diverse applications in digital mapping Hasheminasab et al. (2022), virtual reality Jiang et al. (2020), and smart city Zhou et al. (2020). With a research history spanning decades (Schenk, 1997), classical AAT algorithms can be primarily categorized into two groups: incremental style and global style Schonberger and Frahm (2016). Furthermore, the evolution of deep learning has given rise to numerous supervised AAT algorithms Xiao et al. (2022). The existing AAT methods are reviewed as follows.
1.1 Classic Automated Aerial Triangulation
@@
The first step of the classic AAT is to perform feature extraction and matching of all input images. The following steps are different for the global style and incremental style. Global AAT can predict all camera poses and scene structure at once (Govindu, 2004). In AAT algorithms, Bundle Adjustment (BA) (Triggs et al., 2000) is the most time-consuming part. Global AAT only requires to execute BA once, resulting in higher efficiency. However, it can be difficult to eliminate outliers, resulting in poor robustness and scene integrity. Incremental AAT was first proposed by Snavely et al. (2006), with the key lying in selecting a good initial matching image pair (Beder and Steffen, 2006). Afterward, incremental AAT adds a new image to the system sequentially, followed by Perspective-n-Points (PnP) (Lepetit et al., 2009), Triangulation (Hartley and Sturm, 1997), and local BA. Incremental AAT requires multiple BA, resulting in low reconstruction efficiency in situations with a large number of images (Zhu et al., 2017). In addition, due to the accumulation of errors, the reconstructed scene is prone to drift issues.
@@
Compared to the general scenes, UAV images exhibit distinctive characteristics, including large volumes, high resolutions, and significant overlap. Within the realm of classic AAT algorithms, incremental methods have emerged as the standard approach for UAV image AAT due to their superior robustness against outliers and ability to provide comprehensive results. Addressing the challenges posed by large-scale UAV images, most SOTA AAT methods typically involve employing a divide-and-conquer strategy. This strategy begins by segmenting the UAV image into blocks based on GPS information, followed by the fusion of all modules to yield globally consistent large-scale results. Noteworthy contributions in this field include the work by Chen et al. (2020), which employed the maximum spanning tree to expand images after dividing the scene map into smaller segments with a certain degree of overlap, thereby enhancing connectivity and scene map integrity. Similarly, Xu et al. (2021) introduced a hierarchical approach that constructed a binary tree using images as leaf nodes, subsequently fusing spatial triads and scenes from the bottom up. This method offers advantages in terms of robustness, accuracy, and efficiency. Likewise, Bhowmick et al. (2017) initially organized images into hierarchical trees using clustering methods, then addressed the AAT problem for large-scale images by reconstructing each small image set and merging them into a common reference framework. Snavely et al. (2008) partitioned extensive scenes by calculating the small bone skeleton set and reconstructing the skeleton set of the image. This approach reduces the number of parameters under consideration and enhances reconstruction efficiency. To sum up, global AAT offers high efficiency but suffers from poor robustness and scene integrity. On the other hand, incremental AAT exhibits high robustness and accuracy but tends to have relatively lower time efficiency.
1.2 Supervised Automated Aerial Triangulation
@@
Recognizing the limitations encountered by classical AAT algorithms, an increasing number of studies are exploring the application of deep learning methods to address these challenges. Many existing deep learning methods directly regress the depth map and pose of a monocular camera (Zhou et al., 2017), which usually highly rely on prior information for prediction. In addition, due to not considering the correlation between depth and pose, the generalization ability of these networks is limited, making it difficult to obtain ideal prediction results. BA-Net (Tang and Tan, 2018) attempts to use feature metric BA to solve the AAT problem. It makes end-to-end training possible by designing a differentiable LM (Levenberg–Marquardt) optimization algorithm, but the LM algorithm occupies a large amount of memory and has low computational efficiency. DeepSfM (Wei et al., 2020) can simultaneously regress the pose and depth maps corresponding to the image; however, it requires coarse poses and depth maps for initialization and has high GPU requirements, making it difficult to scale up for high-resolution images and large-scale environments. DeepMLE (Xiao et al., 2022) does not require initial values as input, which expresses the two-view AAT problem as maximum likelihood estimation, learning the relative pose of the two views by maximizing the likelihood function of correlation. Similarly, for the problem of binocular vision estimation, Wang et al. (2021) proposed a dense optical flow estimation network for predicting between two frames, which includes a scale-invariant depth estimation module that can simultaneously calculate the relative pose of the camera based on the corresponding relationship of 2D optical flow. DRO (Gu et al., 2021) is an optimization method based on recurrent neural networks that iteratively updates camera pose and image depth to minimize feature measurement errors. Zhuang and Chandraker (2021) used a self-attention graph neural network to enhance strong interactions between different corresponding relationships and potentially model complex relationships between points to drive learning. MOAC (Wu et al., 2022) introduces a grouped dual cost enhancement module, which enhances the spatial semantic information and channel relationships of costs, making the optimization more robust to noise. Moran et al. (2021) proposed a new approach to solve AAT problems using deep learning. They use matched feature points as input, and after permutation equivariant networks, predict the pose of each camera and 3D points in the scene. Compared to many existing deep learning AAT methods, it can be applied to large-scale reconstruction tasks in an unsupervised manner. However, there are mainly two drawbacks to it. The first one is that it cannot eliminate incorrectly matched point pairs, which means all the input pairs should be correct and is usually difficult to achieve. Another one is that its prediction results are still not satisfactory because of the limited generalizability.
@@
(1) DeepAAT incorporates a spatial-spectral feature aggregation module, specifically combining both the spatial layout and spectral characteristics of an image set. This module boosts the network’s ability to perceive the spatial arrangement of cameras and enhances the global regression capability for poses.
(2) DeepAAT introduces an outlier rejection module according to global consistency, which effectively generates a reliability evaluation score for each feature correspondence. This approach facilitates the efficient and precise elimination of erroneous matching pairs, thereby ensuring accuracy and reliability throughout the entire 3D reconstruction process.
@@
The rest of this paper is structured as follows. The preliminaries for our system are provided in Section 2. A brief system overview including the hardware and software structure is provided in Section 3. A detailed description of DeepAAT is presented in Section 4 and experiments are conducted on UAV image datasets in Section 5. Conclusion and future work are drawn in Section 6.
2 Preliminary
2.1 Problem Definition of Automated Aerial Triangulation
The task of AAT refers to estimating the camera poses and 3D scene points corresponding to the 2D observations on the images. In classic photogrammetry, it is well studied and understood that the relative camera poses and 3D scene points can be solved only with the 2D observations (He and Habib, 2018). The absolute camera poses and 3D scene points respectively to the geodesic framework can be then obtained with Ground Control Points (GCPs) or the GPS mounted on the UAV (Li et al., 2019).
Assume that the stationary targeting survey area is viewed by $M$ images, which are captured by the camera with known pre-calibrated intrinsic parameter $\mathbf{K}$ at different places along the UAV survey mission. The $M$ unknown camera poses are represented by a set of projection matrices $\mathcal{P}=\{\mathbf{P}_{m}|m=1,...,M\}$. Each projection matrix $\mathbf{P}_{m}$ with the size of $3\times 4$ is constructed by camera rotation $\mathbf{R}_{m}\in SO(3)$ (corresponding quaternion is $\mathbf{q}_{m}$) and position $\mathbf{t}_{m}\in\mathbb{R}^{3}$ according to $\mathbf{P}_{m}=[\mathbf{R}_{m}|\mathbf{t}_{m}]$. Given $N$ 3D scene points in the targeting survey area =$\mathcal{F}=\{\mathbf{F}_{n}|n=1,...,N\}$, each 3D scene point is written by $\mathbf{F}_{n}=[\mathbf{F}^{1}_{n},\mathbf{F}^{2}_{n},\mathbf{F}^{3}_{n},1]^{\top}$ in homogeneous coordinates. If $\mathbf{F}_{n}$ can be observed by the $m^{th}$ image, its projection on the $m$’s image is given by Eq.(1). As the depth information $\lambda_{m,n}$ is lost during the projection, $\mathbf{f}_{m,n}$ is an up-to-scale bearing vector.
@@ 2.1 Problem Definition of Automated Aerial Triangulation
In a typical AAT procedure, the initial step involves 2D feature detection and matching between pairwise images, which is carried out using the widely used Scale-Invariant Feature Transform (SIFT) (Lowe, 2004) or other robust feature detectors and descriptors (Dusmanu et al., 2019). This step is not the main focus of this work. Subsequently, a set of 2D feature tracks denoted as $\mathcal{T}=\{\mathbf{T}_{n}|n=1,...,N\}$, is used as the input for our algorithm. It should be noted that track $\mathbf{T}_{n}$ corresponds to the 3D feature $\mathbf{F}_{n}$ and is constructed by a set of 2D observations from different images using Eq.(2):
@@
where $\mathcal{J}_{n}$ denotes the set of images that can observe the 3D feature $\mathbf{F}_{n}$. Then the tracks could be used to recover the camera poses and 3D scene points like the existing incremental (Schonberger and Frahm, 2016) or global (Moulon et al., 2013b) AAT strategies.
2.2 Projective Factorization
Despite the mainstream AAT methods listed above, projective factorization (Sturm and Triggs, 1996) is also a long-established method in AAT. We provide a brief introduction to projective factorization, as it forms the foundation for the operation of the proposed network. The complete image projections, namely the 2D feature tracks $\mathcal{T}$, can be gathered into a measurement matrix $\mathbf{W}_{\mathbf{mes}}$ in Eq.(3):
$\begin{split}\mathbf{W}_{\mathbf{mes}}&\equiv\left[\begin{matrix}\lambda_{1,1}\mathbf{f}_{1,1}&\lambda_{1,2}\mathbf{f}_{1,2}&\cdots&\lambda_{1,N}\mathbf{f}_{1,N}\\ \lambda_{2,1}\mathbf{f}_{2,1}&\lambda_{2,2}\mathbf{f}_{2,2}&\cdots&\lambda_{2,N}\mathbf{f}_{2,N}\\ \vdots&\vdots&\ddots&\vdots\\ \lambda_{M,1}\mathbf{f}_{M,1}&\lambda_{M,2}\mathbf{f}_{M,2}&\cdots&\lambda_{M,N}\mathbf{f}_{M,N}\\ \end{matrix}\right]\\ &=\left[\begin{array}[]{c}\mathbf{KP}_{1}\\ \mathbf{KP}_{2}\\ \vdots\\ \mathbf{KP}_{M}\\ \end{array}\right]\left[\begin{array}[]{c}\mathbf{F}_{1}\\ \mathbf{F}_{2}\\ \vdots\\ \mathbf{F}_{N}\\ \end{array}\right]^{\top}.\end{split}$ (3)
If the 3D scene points in the targeting survey area $\mathcal{F}$ are observed by all the images, the camera poses and 3D scene points can be recovered using Singular Value Decomposition (SVD) of $\mathbf{W}_{\mathbf{mes}}$ (Sturm and Triggs, 1996). For the common cases of missing observations, the SVD can be replaced with iterative methods (Magerand and Del Bue, 2017; Dai et al., 2013). However, these methods are usually too weak for AAT in the presence of outliers and noise (Iglesias et al., 2023), and can not be directly applied to large-scale AAT. Nevertheless, the formulation of $\mathbf{W}_{\mathbf{mes}}$ provides an ideal way of keeping spatial correlation information for neural networks.
2.3 Permutation Equivariant Layer
Let $\mathbf{W}$ be a tensor with the shape of $M\times N\times D$, whose row represents the image index, the column represents the feature track index and the third dimension represents the feature index. Taking measurement matrix $\mathbf{W}_{\mathbf{mes}}$ in Eq.(3) as an example, $\mathbf{W}_{\mathbf{mes}}$ can be rearranged with the shape of $M\times N\times 2$ (the third dimension records the feature coordinates on the image plane) to serve as input for the neural network for the sake of convenience. To recover the camera poses and 3D scene points using a deep neural network, we expect a particular layer can output the same results irrespective of the order of the camera poses or the feature tracks. This reordering problem could be solved using the Permutation Equivariant Layer (PEL)(Hartford et al., 2018), which was first introduced by Moran et al. (2021) to handle SfM problem exploring tensor’s exchangeability.